DETAILS LARTRIG10 3.3.031. Find u + v, u - v, and 2u - 4v. Then sketch each resultant vector. u = (4,3), v = (3,5)
Answer #1

es (7,8) .8 = и X = 4 8 ya us (4,3), V=(3,5) Utv (4,3) + (3,5) = (4+3, 3+5) (7,8) U-V = (4,3) -(3,5) (4-3, 3-5) = (1, -2) au-